OpenStudy (anonymous):

HELP ME PLEASE: solve the system (y=-x^2 - x + 6 (y= x+3

OpenStudy (anonymous):

use substitution

OpenStudy (anonymous):

can quadratic method work

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

OKay then show me what you did since you said yo ucan do the quadratic method

OpenStudy (anonymous):

y=-x^2 - x + 6 (y= x+3 confused

OpenStudy (anonymous):